A boat takes 8 hours to go 6 miles upstream and come back to the starting point. If the speed of the boat is 4 miles per hour in still water, what is the rate of the current?
Answer by stanbon(75887)

A boat takes 8 hours to go 6 miles upstream and come back to the starting point. If the speed of the boat is 4 miles per hour in still water, what is the rate of the current?
-------
Upstream DATA:
time = x hrs ; distance = 6 miles ; rate = d/t = 6/x hrs
----
Downstream DATA:
time = 8-x hrs ; distance = 6 miles ; rate = d/t = 6/(8-x)
--------
Equation:
Upstream Rate Eq::: 4-c = 6/x
Downstream Rt Eq::: 4+c = 6/(8-x)
----------------
Add to solve for "x":
8 = (6/x)+(6/(8-x))
-----
8x(8-x) = 6(8-x) + 6x
-----
64x - 8x^2 = 48
---
x^2 - 8x + 48 = 0
-----
No Real Number solution.
================
Cheers,
Stan H.
================